OpenStudy (anonymous):

Find the points on the graph of (x+6)^2 that are closest to the point (-3,0)

OpenStudy (anonymous):

a point on the graph is \((x, (x+6)^2)\) and the square of the distance will be \[(x+3)^2+(x+6)^4\] minimize that one

OpenStudy (anonymous):

do i have to factor all that out?

OpenStudy (anonymous):

i hope this is a calculus course you have to take the derivative, set it equal to zero and solve to find the critical points

OpenStudy (anonymous):

btw it is already factored, but if you like you could multiply it might make it easier

OpenStudy (anonymous):

yes it is calculus. multiplying to is going to take a long time but thanks! :)

OpenStudy (anonymous):

takes a nano second http://www.wolframalpha.com/input/?i=%28x%2B3%29^2%2B%28x%2B6%29^4

OpenStudy (anonymous):

oh thank you so much! I remember using that website in the past but i forgot the name. Also i have another question. Would you please help me on it: A rectangular solid with a square base has a volume of 6859 cubic inches. (Let x represent the length of the sides of the square base and let y represent the height.) (a) Determine the dimensions that yield the minimum surface area.

OpenStudy (anonymous):

surface area is \(2x^2+4xy\) and since volume is \(6859\) you know \(x^2y=6859\) making \(y=\frac{6859}{x^2}\) and the surface area is now \[S(x)=2x^2+\frac{4\times 6859}{x^3}\]

OpenStudy (anonymous):

ok that was wrong surface area is ow \[S(x)=2x^2+\frac{4\times 6859}{x}\]

OpenStudy (anonymous):

so now you would take the derivative and set it equal to zero right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok thanks! You are a life saver!!! :D

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

ok for x I got 19 but how do you find the other side? do you plug it back into the original?

OpenStudy (anonymous):

oh nevermind i got it